A certain grade of coal contains 1.7 % S. Assume the burning of S compounds can be presented by the following equation:


                             S + O2 → SO2   



Calculate the amount of the pollutant SO2 gases emitted in the atmosphere, per ton (1,000 kg,) of coal burned if the process is 79% efficient.



1
Expert's answer
2021-01-19T04:30:06-0500

At first, the mass of S in 1 ton of coal may be calculated:

m(S)=1000×0.017=17(kg)=17000(g).

Theoretically, according to the chemical equation, from one mol of S one mol of SO2 is obtained:

n=m/M

n(SO2)=n(S)=17000/32=531.25(mol).

On practice:

n(SO2)=531.25×0.79=419.6875(mol).

m=n×M

M(SO2)=32+2×16=64(g/mol)

m(SO2)=419.6875×64=26860(g)=26.86(kg).